OpenStudy (anonymous):

Find the Limit: lim x-0 (sinsxsin5x)/(x^2)

ganeshie8 (ganeshie8):

familiar with the limit : \[\large \lim\limits_{x\to 0} \dfrac{\sin x}{x} = 1\] ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

use it

OpenStudy (anonymous):

how? can I pull a sinx out of the top? making it sinx(15)/x^2? then i could make it (sinx/x)*(15/X)?

ganeshie8 (ganeshie8):

\[\large \lim\limits_{x\to 0} \dfrac{\sin 3x~\sin 5x}{x^2}\]

ganeshie8 (ganeshie8):

is that the limit you're trying to evaluate ?

OpenStudy (anonymous):

yes! I just realized I had a typo in the original

ganeshie8 (ganeshie8):

\[\large \begin{align}\lim\limits_{x\to 0} \dfrac{\sin 3x~\sin 5x}{x^2} &=15~\lim\limits_{x\to 0} \dfrac{\sin 3x}{3x}\dfrac{\sin 5x}{5x} \end{align} \]

ganeshie8 (ganeshie8):

split that into product of two limits each evaluates to 1

ganeshie8 (ganeshie8):

\[\large \begin{align}\lim\limits_{x\to 0} \dfrac{\sin 3x~\sin 5x}{x^2} &=15~\lim\limits_{x\to 0} \dfrac{\sin 3x}{3x}\dfrac{\sin 5x}{5x} \\~\\&= 15~\left( \lim\limits_{3x\to 0} \dfrac{\sin 3x}{3x}\right) ~ \left(\lim\limits_{5x\to 0} \dfrac{\sin 5x}{5x} \right)\\~\\ &= 15~\left( 1\right) ~ \left(1\right)\\~\\&= 15\end{align} \]

OpenStudy (anonymous):

I forgot you can split the limits like that and multiply a number through like that... thank you so much! I didn't even think about looking back at my limit rules.. =(

ganeshie8 (ganeshie8):

we cannot do that everytime, but we can do it whenever we know that both the limits exist

ganeshie8 (ganeshie8):

\[\large \lim\limits_{x\to a} f(x)\times g(x) = \lim\limits_{x\to a} f(x)\times \lim\limits_{x\to a} g(x) \] is true only when both the limits on right hand side exist

ganeshie8 (ganeshie8):

it may not look like a big deal, but it can get us into weird problems if we are not careful..

OpenStudy (anonymous):

so if the limit doesn't exist you cannot split them because it would be like multiplying by something that isn't there?

ganeshie8 (ganeshie8):

Exactly !

OpenStudy (anonymous):

sorry I had to put that into simple english. Thank you so much! I know I am new but I love algebra and as a user am I allowed to go help others?

ganeshie8 (ganeshie8):

definitely ! you seem new around here ? let me take a moment to welcome you officially :) \[\large \color{green}{\text{Welcme to Openstudy !!}}\]

ganeshie8 (ganeshie8):

when you're free, click below to know more about how this site works and other stuff http://openstudy.com/code-of-conduct have fun !

OpenStudy (anonymous):

Thanks!